OpenStudy (anonymous):

1.whats negative 3 and 4/5 as an improper fraction 2.whats 6 and 5/8 as a improper fraction 3. What's 2 and 2/9 as an improper fraction

OpenStudy (anonymous):

1: 19/5

OpenStudy (anonymous):

2: 53/8

OpenStudy (anonymous):

3: 20/9

OpenStudy (anonymous):

thats the answers

OpenStudy (lizz123):

\[-3\frac{ 4 }{ 5 }\] Multiply 5 and -3 15+4= 19 19 has to go over the denominator which in this case is 5 So you answer will be \[\frac{ 19 }{5 }\]

OpenStudy (lizz123):

In question 2 \[6\frac{5 }{8 }\] Multiply the denominator is 8 so you multiply the denominator and the whole number \[8\times6=48\] Now add 48 to the number: \[48+5=53\] Now the numerator is 53 and the denominator is the same So you're answer is \[\frac{ 53 }{ 8}\]